Diana has available 400 yards of fencing and wishes to enclose a rectangular area.
​(a) Express the area A of the rectangle as a function of the width W of the rectangle.
​(b) For what value of W is the area​ largest?
​(c) What is the maximum​ area?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Draw this
sides are x x and 400-2x
The area is x(400-2x)=-2x^2+400x
the vertex for this quadratic is a maximum and will have x value of -b/2a=-400/-4=100 yds
the maximum area for this 100*200 rectangle is 20000 yds^2.
The formula for A=-2W2+400W for W=width

Answer by ikleyn(52830)

.

            The post by  @boreal  represents the solution in a very inaccurately way,  and is incorrect in many places.

            I came to bring the correct solution.


sides are x and 200-x    (where 200 = 400/2  is half of the perimeter)


The area is x(200-x) = -x^2 + 200x


the vertex for this quadratic is a maximum and will have x value of 


    -b/2a = (-200)/(-2) = 100 yds


The maximum area for this 100*100   SQUARE (!)  is 10000 yds^2.


The formula for the area is   = -W^2 + 200*W  for  W= width.

----------------

The answer is very well known fact:  under given condition,  the maximum area is provided by the square,
and when the perimeter of a rectangle is given  (= the fence length),  the side of this square is one fourth of the perimeter.


See the lesson
    - A rectangle with a given perimeter which has the maximal area is a square
    - A farmer planning to fence a rectangular garden to enclose the maximal area
in this site.

On finding the maximum/minimum of a quadratic function see the lessons
    - HOW TO complete the square to find the minimum/maximum of a quadratic function
    - Briefly on finding the minimum/maximum of a quadratic function
    - HOW TO complete the square to find the vertex of a parabola
    - Briefly on finding the vertex of a parabola

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Finding minimum/maximum of quadratic functions".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.